OpenStudy (karlaltr):

Tia is investing $ 2500 that she would like to grow to $ 6000 in 10 years, at what annual interest rate compounded quarterly must Tia invest her money? I know that PV = 6000 A= 2500 i = ? n = 40 My teacher gave us the formula PV = A (1+i) -n I am having troubles isolating i, I am not sure how to get to the right answer

OpenStudy (welshfella):

use the compound interest formula A = P(a + r/t) ^ nt where P is the starting amount , r = annual rate in decimal form, t = number of times interest is pais in a year and n = number of years

OpenStudy (karlaltr):

So how would I get rid of thee exponent? square root?

OpenStudy (welshfella):

plug in the given values first 6000 = 2500( 1 + r/4)^40

OpenStudy (karlaltr):

Im confused as to why it is r/4

OpenStudy (karlaltr):

is it because i = r/N

OpenStudy (welshfella):

because the interest is paid 4 times a year r = annual rate

OpenStudy (welshfella):

That is what they ask for - annual rate.

OpenStudy (karlaltr):

alright that makes sense

OpenStudy (welshfella):

so if we divide both sides by 2500 to isolate the exponent part we get (1 + r/4)^40 = 2.4 now take logs of both sides

OpenStudy (karlaltr):

why do we log?

OpenStudy (karlaltr):

Sorry I just want to make sure I understand this well so I can do it again later

OpenStudy (welshfella):

- that gets rid of the exponenents

OpenStudy (karlaltr):

alright so I log the other side

OpenStudy (karlaltr):

the side witht he 2.4

OpenStudy (karlaltr):

is this inverse log or just log?

OpenStudy (welshfella):

what? I somehow lost my post!!???

OpenStudy (welshfella):

take logs - u can use natural logs:- ln (1 + r/4)^40 = ln 2.4 by law of logs 40 ln (1 + r/4) = ln2.4 ln ( 1 + r/4) = ln2.4 / 40 ok so far?

OpenStudy (karlaltr):

I think so, so by doing log I will be getting rid of the 40 exponent?

OpenStudy (welshfella):

yea

OpenStudy (karlaltr):

Alright that makes sense

OpenStudy (karlaltr):

so I have 0.380211242 = 1+ r/4

OpenStudy (welshfella):

now work out right side ln (1 + r/4) = 0.0218867

OpenStudy (karlaltr):

Wait how is it 0.021, Im not sure why Im getting a diffrerent number

OpenStudy (welshfella):

0.021 = ln (1 + r/4) now you can use inverse logs

OpenStudy (welshfella):

you should use 0.0218867 to get accurate result

OpenStudy (welshfella):

use the e^x key on ur calculator - remember we are using natural logs

OpenStudy (karlaltr):

Alright I think I understand now

OpenStudy (welshfella):

( 1 + r/4) = e^0.021887

OpenStudy (welshfella):

good you should get a result which is between 8 and 9%

OpenStudy (welshfella):

Aim to isolate the expression you want to evaluate and then takes logs note we use the log law log a^n = n log a

OpenStudy (welshfella):

It's nice to deal with someone who wants the learn rather than just get an answer.

OpenStudy (karlaltr):

Alright I think that should be alright then, thank you very very much!!

OpenStudy (welshfella):

yw